OpenStudy (anonymous):

How would you simplify (49/27)^(3/2)? Explain please, thanks!

hartnn (hartnn):

NOTE : 49=7^2 27= 3*9 = 3*3^2 ok?

OpenStudy (anonymous):

.. what? /:

OpenStudy (anonymous):

Please read before asking "... what? /:"

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

Where do you don't understand?

hartnn (hartnn):

ok, actually 49 is 7*7 so i can write 49 as \(7^2\) this makes sense?

OpenStudy (anonymous):

Yees.

hartnn (hartnn):

similarly, 27 = 3*9 and 9=3*3 so i wrote 9 as \(3^2\) and 27 = 3*\(3^2\) ok?

OpenStudy (anonymous):

Mhm o:

hartnn (hartnn):

is that a YES ?

OpenStudy (anonymous):

yes.

hartnn (hartnn):

ok, so now i write 49/27 = 7^2 / (3*3^2) and (49/27)^(3/2) = (7^2/(3*3^2))^(3/2)

OpenStudy (anonymous):

Okay.

hartnn (hartnn):

 understood this next step ? i just took 2 inside. using \((a^n)^{1/2}=a^{(n/2)}\)